OpenStudy (anonymous):

Tell whether the lines for the pair of equations are parallel, perpendicular, or neither y=-4/5x+3 4x-5y=-15

OpenStudy (anonymous):

Find the slopes of these lines. If they're identical, they're parallel. If they are negative reciprocals, they're perpendicular. By negative reciprocals, I mean -3 and 1/3 as an example.

OpenStudy (anonymous):

you need to find the slopes of the lines y = -4/5x + 3 --- here the slope is -4/5 4x - 5y = -15 -5y = -4x - 15 y = 4/5x + 3 -- slope here is 4/5 The slopes are not the same, so they are not parallel. The slopes are not negative reciprocals of each other, so they are not perpendicular. That only leaves one....neither

OpenStudy (anonymous):

Thank you tons!